# SyncTwin: Fast Digital Twin Construction and Synchronization for Safe Robotic Grasping

### v1

@@ 3.3 Online Digital Twin Synchronization
The second stage of our system focuses on real-time object tracking and safe grasp execution through continuous perception, planning synchronization between the real and digital environments. This stage consists of four tightly integrated modules: real-time point cloud segmentation, GPU-accelerated colored-ICP registration, grasp pose generation from complete object models, and dynamic motion planning with cuRobo MPC.
@@
Real-time Point Clouds Segmentation. To achieve real-time segmentation on incoming RGB-D streams, we build a module shown in Fig 4, that performs continuous inference on camera frames and outputs segmentation masks $\mathcal{S}_{p}$, and then projects the mask onto the full point cloud to obtain the corresponding partial object point clouds $\mathcal{X}_{p}$. Compared with traditional offline SAM2 inference, we design a sliding window mechanism that enables SAM2 to process camera streams in real time, maintaining temporal consistency in object masks for continuous 3D segmentation and tracking under occlusions. At image size 640×480, our method runs at 15 Hz on RTX 4090.
Colored-ICP Registration. For aligning partial point clouds $\mathcal{X}_{p}$ obtained from the camera with the corresponding full object model $\mathcal{X}_{m}$ stored in the memory bank, we employ a colored-ICP algorithm [32] implemented on the GPU via the cupoch library [42]. Unlike traditional geometric ICP, which minimizes only spatial distance, colored-ICP jointly minimizes geometric and color residuals: $E(R,t)=\sum_{i}[\lambda_{g}\|R\mathbf{x}_{i}+t-\mathbf{y}_{i}\|^{2}+\lambda_{c}\|I(\mathbf{x}_{i})-I(\mathbf{y}_{i})\|^{2}]$ where $(R,t)$ denotes the transformation between $\mathcal{X}_{p}$ and $\mathcal{X}_{m}$, and $I(\cdot)$ represents point color intensity. The weighting factors $\lambda_{g}$ and $\lambda_{c}$ balance geometric and color terms.
@@
Figure 4: Overview of the camera predictor module. The red solid line indicates that the first frame is persistently stored in the frame memory. The yellow dashed line represents the operation of saving frames from the previous time step into the memory. The blue and green solid lines denote the data flow and processing steps for the current frame. Together, the sliding-window mechanism enables real-time video segmentation and object-level point cloud tracking with temporally consistent memory updates.
@@
Grasp Pose Generation and Obstacle Representation. For unseen target objects, we directly apply GraspGen to the partial point cloud. In contrast, once a seen target object is registered, we replace its partial observation with the complete point cloud $\mathcal{X}_{m}$ from the memory bank and feed it into GraspGen to predict grasp poses $\{\mathbf{T_{gripper}}=f_{\text{GraspGen}}(\mathcal{X}_{m})\}$,This replacement mitigates the uncertainty from occlusion and single-view perception, yielding more stable and accurate grasp pose estimation. For unseen obstacles, we dynamically generate multi-convex hulls (V-HACD [49]) from the segmented point cloud for collision modeling. For known obstacles, the aligned object meshes and poses are directly imported into the digital twin for real-time collision checking.
@@
Motion Planning with Sim-to-real Synchronization. Finally, motion planning and control are performed using the cuRobo’s model predictive control (MPC) framework. At each control step, the robot’s joint states are synchronized with the digital twin, where cuRobo computes an optimized short-horizon trajectory under real-time collision constraints. Only the first control action $\{\mathbf{a}_{0}\}$ from the predicted trajectory $\mathbf{A}_{t:t+H}=\{\mathbf{a}_{0},\mathbf{a}_{1},\ldots,\mathbf{a}_{H}\}$ is executed on the real robot, followed by continuous replanning with the environment updates, achieving closed-loop synchronization between simulation and reality.
4 Experiments
@@
We evaluate SyncTwin across three dimensions: (1) the efficiency of the fast 3D reconstruction pipeline; (2) obstacle avoidance performance under dynamic and single-view occluded conditions; and (3) grasp success rate under single-view occlusion. All experiments are designed to validate both the offline and online components of our framework.
4.1 Experiment Setup
@@
All experiments are conducted with a Franka Emika Panda robotic arm equipped with an Intel RealSense D455 RGB-D camera mounted above and in front of the workspace. And iPhone 12 for RGB images. We apply voxel-based downsampling to the input point cloud with a voxel size of 3 mm. This is motivated by the fact that the computational complexity of point cloud segmentation algorithms scales linearly with the number of points. The digital twin is implemented in Isaac Sim 4.0 and integrated with the cuRobo MPC motion planning framework, running on a single NVIDIA RTX 4090 GPU. Both perception and planning run on the same GPU, enabling a closed-loop update rate of up to 5 Hz. Motion planning runs at 10 Hz, and the robot’s velocity scaling is set to 0.2. The test objects include bottles, cans, cups, and boxes of various shapes.
@@
Figure 5: The comparison of processing time given different numbers of input images. The processing time covers both reconstruction and segmentation. The 5 and 10 images do not apply to the baselines because of fail to estimate the camera extrinsic.
4.2 Baselines and Metrics
@@
Baselines. (1) For 3D reconstruction: we compare against Photogrammetry [35], NeRF (Nerfstudio [43]), and Gaussian Splatting (3DGS [20]). (2) For obstacle avoidance: we adopt NVBlox [23] as the baseline voxel-mapping method. For the ablation studies, we analyze performance across the following settings: (1) Mask Expansion and Denoising: using variants without mask expansion and without denoising as baselines, evaluating their reconstruction quality against our full segmentation mechanism. (2) Object Completion: we use GraspGen [26] as the baseline grasp generator, comparing grasp poses predicted from single-view partial point clouds (baseline) versus from complete, asset-retrieved geometry produced by SyncTwin (ours).
@@
Evaluation Metrics. Depending on the experiment, we report the following metrics: reconstruction time (min), dependency on the number of input images $N_{\min}$, obstacle avoidance success rate (%), and grasp success rate (%). For avoidance tests, we define 3-level outcome levels: FA (full avoidance, no contact with the object, 1.0), EA (edge avoidance, slight contact without displacing the object, 0.8), and CO (collision, noticeable contact that significantly moves the object, 0.0). Weighted success rate:
@@
Table 1: Comparison of obstacle-avoidance performance between NVBlox and SyncTwin in dynamic environments. Unseen objects (left) are not present in the asset memory, whereas starred objects in the Seen category (right) are stored in the memory bank. SyncTwin achieves significantly higher success rates in both settings, with particularly strong gains when complete object assets are available.
@@
4.3 Fast 3D Reconstruction
We evaluate the efficiency and input-image dependency of SyncTwin’s 3D reconstruction module, comparing against Photogrammetry, 3DGS, and Nerfstudio. Each method reconstructs the same object using 5, 10 ,20, 40, 60 RGB images, and all approaches obtain object assets through multi-view projection based segmentation. The experimental results are summarized as follows.
Reconstruction Time. As shown in Figure 5, SyncTwin achieves the shortest reconstruction time across all settings. It produces a simulation-ready mesh in only about 1–2 minutes using 5–10 input images, while Photogrammetry, 3DGS, and Nerfstudio require at least 4–7 minutes even with more frames. This enables significantly faster digital-twin construction in a new real-world scenarios.
Dependency on the Number of Input Images. SyncTwin also exhibits the lowest dependency on input-image count. It generates usable meshes from as few as 5–10 images, whereas competing approaches typically require 20+ images to avoid failure caused by unstable optimization of camera extrinsics. This low image dependency substantially accelerates asset generation and improves robustness under limited views, especially because the reported processing time excludes image-capture time, which becomes additional and unpredictable when more images are required.
Furthermore, the qualitative comparison in Figure 6 shows that our method preserves fine-grained geometric details even with very few input images, as our mesh simplification algorithm maintains high geometric fidelity while reducing mesh vertices, enabling faster simulation performance. For instance, the bear’s ear shape remains well preserved with only 5–10 frames, demonstrating both low image dependency and strong geometric consistency.
These properties make SyncTwin a state-of-the-art solution for fast 3D geometry asset generation: the system can construct high-quality, simulation-ready object meshes in about 1 minute, enabling rapid reconstruction of object assets for downstream digital-twin synchronization.
4.4 Obstacle Avoidance under Occlusion
@@
Both NVBlox and SyncTwin are tested on unseen (not stored in memory bank) obstacles to evaluate avoidance performance. Additionally, SyncTwin is evaluated on seen (stored in memory bank) objects recorded in its asset memory to study how prior geometry improves safety and reactivity under single-view occlusion. The experiment tests under two motion patterns: SelfRot indicates in-place rotation; EnterTraj blocks motion into the predicted trajectory. Each condition is repeated for 20 trials, total $N=20\times 4=80$. We test four boxes with sizes of Box1: $10{\times}10{\times}10$ cm, Box2: $20{\times}20{\times}20$ cm, Box3: $10{\times}20{\times}30$ cm, and Box4: $20{\times}22{\times}35$ cm as obstacles. Through experimental observation, the results are as follows:
Unseen Object Performance. As shown in Table 1, SyncTwin consistently outperforms NVBlox when encountering unseen obstacles under single-view dynamic occlusion. NVBlox exhibits failure modes such as: (1) unstable voxelization for small objects (e.g., Box1) due to sparse depth returns; (2) misclassification of limited height objects (e.g., Box3) as part of the tabletop caused by inconsistent depth estimation; and (3) trajectory intersections when large obstacles (e.g., Box4) fall outside the sensor’s visible region. Across all these cases, SyncTwin maintains markedly higher obstacle-avoidance success rates, demonstrating stronger robustness to object scale, occlusion, and limited viewpoint coverage.
Seen Object Performance. In Table 1, when objects are stored in the asset memory, SyncTwin’s performance improves even further. Complete geometric priors obtained in Stage I resolve the challenges posed by small or thin objects, eliminate many failure modes, and convert numerous edge-avoidance cases into full avoidance. These results indicate that SyncTwin not only generalizes better on unseen objects but also achieves substantially higher reliability when the objects have been previously reconstructed. These results highlight the importance of SyncTwin’s Stage I memory construction, which enables complete geometric reasoning even under partial observability. Furthermore, Figure 7 illustrates representative avoidance examples achieved by SyncTwin.
4.5 Ablation Studies
Mask Expansion and Denoising. Due to the inaccuracy of VGGT-estimated camera extrinsics, multi-view mask projections often become misaligned and fail to fully cover the object. Our ablation study evaluates reconstruction results w/o mask expansion and w/o denoising. As shown in Figure 8, removing margin expansion causes inconsistent projections across viewpoints, leaving parts of the object missing. Likewise, disabling supporting-plane points denoising preserves table artifacts in the point cloud, resulting in noisy meshes. These comparisons demonstrate that segmentation-aware expansion and support-plane denoising are essential for obtaining clean and stable 3D assets.

### v2

@@ III-C Online Digital Twin Synchronization
The second stage of our system focuses on real-time object tracking and safe grasp execution through continuous perception, planning synchronization between the real and digital environments. This stage consists of four integrated modules: real-time point cloud segmentation, GPU-accelerated colored-ICP registration, grasp pose generation from complete object models, and dynamic motion planning with cuRobo MPC.
@@
Real-time Point Clouds Segmentation. To achieve real-time segmentation on incoming RGB-D streams, we build a module shown in Fig. 4, that performs continuous inference on camera frames and outputs segmentation masks $\mathcal{S}_{p}$, and then projects the mask onto the full point cloud to obtain the corresponding partial object point clouds $\mathcal{X}_{p}$. To support streaming input, we design a sliding-window inference mechanism that maintains temporal consistency of object masks across frames, enabling continuous 3D segmentation and tracking under occlusions. At image size 640×480, our method runs at 15 Hz on RTX 4090.
@@
Figure 4: Overview of the camera predictor module. The red solid line indicates that the first frame is persistently stored in the frame memory. The yellow dashed line represents saving frames from the previous time step into memory. The blue and green solid lines denote the data flow and processing of the current frame. Together, the sliding-window mechanism enables real-time video segmentation and object-level point cloud tracking with temporally consistent memory updates.
Colored-ICP Registration. For aligning partial point clouds $\mathcal{X}_{p}$ obtained from the camera with the corresponding full object model $\mathcal{X}_{m}$ stored in the memory bank, we employ a colored-ICP algorithm [55] implemented on the GPU via the cupoch library [9]. Unlike traditional geometric ICP, colored-ICP jointly minimizes geometric and color residuals: $E(R,t)=\sum_{i}\left[\lambda_{g}\|R\mathbf{x}_{i}+t-\mathbf{y}_{i}\|^{2}+\lambda_{c}\|I(\mathbf{x}_{i})-I(\mathbf{y}_{i})\|^{2}\right]$. Here, $(R,t)$ denotes the rigid transformation between $\mathcal{X}_{p}$ and $\mathcal{X}_{m}$, and $I(\cdot)$ represents point color intensity. The weighting factors $\lambda_{g}$ and $\lambda_{c}$ balance geometric and color terms.
@@
Grasp Pose Generation and Obstacle Representation. For unseen target objects, we directly apply GraspGen to the partial point cloud. In contrast, once a seen target object is registered, we replace its partial observation with the complete point cloud $\mathcal{X}_{m}$ from the memory bank and feed it into GraspGen to predict grasp poses $\{\mathbf{T_{gripper}}=f_{\text{GraspGen}}(\mathcal{X}_{m})\}$. This replacement mitigates the uncertainty from occlusion and single-view perception, yielding more stable and accurate grasp pose estimation. For unseen obstacles, we dynamically generate multi-convex hulls (V-HACD [56]) from the segmented point cloud for collision modeling. For known obstacles, the aligned object meshes and poses are directly imported into the digital twin for real-time collision checking.
@@
Motion Planning with Real2sim2real Synchronization. Finally, motion planning and control are performed using cuRobo’s MPC framework. At each control step, the robot’s joint states are synchronized with the digital twin, where cuRobo computes an optimized short-horizon trajectory under real-time collision constraints. Only the first control action $\{\mathbf{a}_{0}\}$ from the predicted trajectory $\mathbf{A}_{t:t+H}=\{\mathbf{a}_{0},\mathbf{a}_{1},\ldots,\mathbf{a}_{H}\}$ is executed on the real robot, followed by continuous replanning with environment updates, achieving closed-loop synchronization between simulation and reality.
IV Experiments
@@
We evaluate SyncTwin across four dimensions: (1) the efficiency of the fast 3D reconstruction pipeline; (2) the pose accuracy of the synchronized digital twin under both unoccluded and occluded conditions; (3) obstacle avoidance performance under dynamic and single-view occluded conditions; and (4) grasp success rate under single-view occlusion. All experiments are designed to systematically validate both the offline and online components of our framework.
IV-A Experiment Setup
@@
All experiments are conducted with a Franka Panda robotic arm equipped with an Intel RealSense D455 RGB-D camera mounted above and in front of the workspace. And iPhone 12 for RGB images. We apply voxel-based downsampling to the input point cloud with a voxel size of 3 mm. This is motivated by the fact that the computational complexity of point cloud segmentation algorithms scales linearly with the number of points. The digital twin is implemented in Isaac Sim 4.0 and integrated with the cuRobo MPC motion planning framework, running on a single NVIDIA RTX 4090 GPU. Both perception and planning run on the same GPU, enabling a closed-loop update rate of up to 5 Hz. Motion planning runs at 10 Hz, and the robot’s velocity scaling is set to 0.2. The test objects include bottles, cans, cups, and boxes of various shapes.
@@
Figure 5: Comparison of processing time given different numbers of input images. The reported time includes both reconstruction and segmentation. In practice, our method achieves reliable reconstruction with as few as 5-10 input images. Results for 5 and 10 images are not reported for the baselines, as they fail to estimate camera extrinsics.
IV-B Baselines and Metrics
@@
Baselines. (1) For 3D reconstruction: we compare against Photogrammetry [17], NeRF (Nerfstudio [57]), and Gaussian Splatting (3DGS [25]); (2) For obstacle avoidance: we adopt NVBlox [5] as the baseline voxel-mapping method. Ablation Studies. We analyze performance across the following settings: (1) Mask Expansion and Denoising: using variants without mask expansion and without denoising as baselines, evaluating their reconstruction quality against our full segmentation mechanism; (2) Object Completion: we use GraspGen [58] as the baseline grasp generator, comparing grasp poses predicted from single-view partial point clouds (baseline) versus from complete geometry produced by SyncTwin (ours).
@@
Evaluation Metrics. Depending on the experiment, we report the following metrics: reconstruction time (min), dependency on the number of input images $N_{\min}$, pose accuracy (%), obstacle avoidance rate (%), and grasp success rate (%).
Pose accuracy is evaluated using normalized translation and rotation accuracy scores. For translation, let $\Delta\mathbf{t}_{\mathrm{real}}^{(i)}$ and $\Delta\mathbf{t}_{\mathrm{twin}}^{(i)}$ denote the object translations in the real world and in the digital twin for the $i$-th trial. The translation error is defined as $e_{\mathrm{trans}}^{(i)}=\left\|\Delta\mathbf{t}_{\mathrm{real}}^{(i)}-\Delta\mathbf{t}_{\mathrm{twin}}^{(i)}\right\|_{2},$ and the translation accuracy is computed as $\mathrm{Acc}_{\mathrm{trans}}=1-\frac{\overline{e}_{\mathrm{trans}}}{t_{\mathrm{real}}}$. For rotation, object orientations are represented as unit quaternions. Given the real and simulated orientations $\mathbf{q}_{\mathrm{real}}^{(i)}$ and $\mathbf{q}_{\mathrm{twin}}^{(i)}$, the angular rotation error is computed as $e_{\mathrm{rot}}^{(i)}=2\arccos\!\left(\left|q_{\mathrm{rel}}^{w}\right|\right)$, where $\mathbf{q}_{\mathrm{rel}}^{(i)}=\left(\mathbf{q}_{\mathrm{real}}^{(i)}\right)^{-1}\otimes\mathbf{q}_{\mathrm{twin}}^{(i)}$. The rotation accuracy is defined as $\mathrm{Acc}_{\mathrm{rot}}=1-\frac{\overline{e}_{\mathrm{rot}}}{r_{\mathrm{real}}}$. Higher accuracy values indicate better alignment between the digital twin and the real world.
For avoidance tests, we define 3-level outcome levels: FA (full avoidance, no contact with the object, 1.0), EA (edge avoidance, slight contact without displacing the object, 0.8), and CO (collision, noticeable contact that significantly moves the object, 0.0). Weighted success rate:
@@
IV-C Fast 3D Reconstruction
We evaluate the efficiency and input-image number dependency of SyncTwin’s 3D reconstruction module, comparing against Photogrammetry, 3DGS, and Nerfstudio. Each method reconstructs the same object using 5, 10, 20, 40, and 60 RGB images as input, and all approaches obtain object assets through multi-view projection-based segmentation. The experimental results are summarized as follows for comparison.
Reconstruction Time. As shown in Fig. 5, SyncTwin achieves the shortest reconstruction time across all settings. It generates a simulation-ready mesh in approximately 1-2 minutes using only 5-10 input images, whereas Photogrammetry, 3DGS, and Nerfstudio require more input images and substantially longer processing times, typically ranging from 4 to 8 minutes. As a result, SyncTwin enables significantly faster digital twin construction in new real-world scenarios.
Dependency on the Number of Input Images. SyncTwin also exhibits the lowest dependency on input-image count. It can generate usable meshes from as few as 5-10 images, whereas competing methods typically require more than 20 images to avoid reconstruction failures caused by unstable camera extrinsic optimization. This low image dependency substantially accelerates asset generation and improves robustness under limited viewpoints, especially because the reported processing time excludes image-capture time, which becomes additional and unpredictable when more images are required.
Figure 7: Illustration of the pose accuracy evaluation setup. The left column shows the real-world scene, where the target object (highlighted in blue) is tracked and segmented during object translation, rotation, and occlusion. The right column visualizes the corresponding synchronized digital twin in simulation after object matching and pose alignment.
TABLE I: Pose accuracy of SyncTwin. Translation (Trans.) and rotation (Rot.) accuracy are evaluated on different objects under unoccluded and occluded conditions.
Box Cup Bottle Ketchup
Condition Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot.
Unoccluded (%) 95.1 78.3 96.3 88.7 94.8 90.4 94.7 82.0
Occluded (%) 85.2 70.4 87.7 78.2 90.3 82.6 91.5 76.3
Furthermore, the qualitative comparison in Fig. 6 shows that our method preserves fine-grained geometric details even with very few input images, as our mesh simplification algorithm effectively maintains high geometric fidelity while reducing the number of mesh vertices, enabling faster simulation performance. For instance, the bear’s ear shape remains well preserved with only 5-10 frames, demonstrating both low image dependency and strong geometric consistency.
These properties make SyncTwin a state-of-the-art solution for fast 3D geometry asset generation: the system can construct high-quality, simulation-ready object meshes in about 1 minute, enabling rapid reconstruction of object assets for downstream digital twin synchronization and safe geometry-aware motion planning in real-world environments.
TABLE II: Comparison of obstacle-avoidance performance between NVBlox and SyncTwin in dynamic environments. Unseen objects (left) are not present in the asset memory, whereas starred objects in the Seen category (right) are stored in the memory bank. SyncTwin achieves significantly higher success rates in both settings, with particularly strong gains when complete object assets are available. Results are reported using the evaluation metrics defined in Sec. IV-B.
@@ IV-C Fast 3D Reconstruction
IV-D Pose Accuracy
We conduct an experiment to evaluate the pose accuracy of the synchronized digital twin. Four object categories (box, cup, bottle, and ketchup bottle) are tested under both unoccluded and single-view occluded conditions. Each object is translated by 15 cm along the Cartesian axes and rotated by 90 degrees around the roll, pitch, and yaw axes, with each motion repeated 50 times, as shown in Fig. 7. As shown in Table I, SyncTwin achieves high translation and rotation accuracy across all object categories in the unoccluded setting, with translation accuracy consistently above 90%. Under occlusion, accuracy decreases but remains stable across objects, demonstrating robust pose synchronization under partial observations.
IV-E Obstacle Avoidance under Occlusion
@@
Both NVBlox and SyncTwin are tested on unseen (not stored in the memory bank) obstacles to evaluate avoidance performance. Additionally, SyncTwin is further evaluated on seen (stored in memory bank) objects recorded in its asset memory to study how prior geometry improves safety and reactivity under single-view occlusion. The experiment tests under two motion patterns: SelfRot indicates in-place rotation; EnterTraj blocks motion into the predicted trajectory. Each condition is repeated for 20 trials, in total $N=20\times 4=80$. We test four boxes with sizes of Box1: $10{\times}10{\times}10$ cm, Box2: $20{\times}20{\times}20$ cm, Box3: $10{\times}20{\times}30$ cm, and Box4: $20{\times}22{\times}35$ cm as obstacles. Based on these experimental settings, the results are summarized as follows:
Unseen Object Performance. As shown in Table II, SyncTwin consistently outperforms NVBlox when encountering unseen obstacles under single-view dynamic occlusion. NVBlox exhibits failure modes such as: (1) unstable voxelization for small objects (e.g., Box1) due to sparse depth returns; (2) misclassification of limited height objects (e.g., Box3) as part of the tabletop caused by inconsistent depth estimation; and (3) trajectory intersections when large obstacles (e.g., Box4) fall outside the sensor’s visible region. Across all these cases, SyncTwin maintains markedly higher obstacle-avoidance success rates, demonstrating stronger robustness to object scale, occlusion, and limited viewpoint coverage.
Seen Object Performance. As shown in Table II, SyncTwin achieves further performance improvements when objects are stored in the asset memory. Complete geometric priors obtained in Stage I resolve the challenges posed by small or thin objects, eliminate many failure cases, and convert numerous edge-avoidance cases into full avoidance. These results indicate that SyncTwin not only generalizes better on unseen objects but also achieves substantially higher reliability when the objects have been previously reconstructed, highlighting the importance of SyncTwin’s Stage I memory construction, which enables complete geometric reasoning even under partial observability. Furthermore, Fig. 8 illustrates representative avoidance examples achieved by SyncTwin.
IV-F Ablation Studies
Mask Expansion and Denoising. Due to inaccurate VGGT-estimated camera extrinsics, multi-view mask projections often become misaligned and fail to fully cover the object. Our ablation study evaluates reconstruction results w/o mask expansion and w/o denoising. As shown in Fig. 9, removing margin expansion causes inconsistent projections across viewpoints, leaving parts of the object missing. Likewise, disabling supporting-plane points denoising preserves table artifacts in the point cloud, resulting in noisy meshes. These comparisons demonstrate that mask expansion and support-plane denoising are essential for obtaining clean 3D assets.